OpenStudy (anonymous):

Replace the ? with a polynomial to make a true statement: 2x -2 + x -1-3 = x -2 (?)

OpenStudy (anonymous):

so get your like terms together on the left. 2x-x=-1x -2-1-3=-6 so you will have -1x-6=x-2 add something that will make both sides equal. I think this is how you do it.